# Selective Prompt Anchoring for Code Generation

### v4

@@ Appendix D Code Generation Examples
In the first example, CodeLlama (7B) overlooks the specified condition "upper vowels." In contrast, Spa enhances the model’s focus on the intended purpose. The code initializes all the upper vowels in the first line and correctly refers to it later.
@@ Appendix E Computational Cost of Spa
We optimize Spa by KV-cache [33] and flash attention [9] to speed up inference. We notice when attention augmentation in Spa is activated, the additional overhead is around 0.6 times the base model’s latency. However, the attention augmentation does not activate for failed tasks. The more accurate the base model, the less overhead Spa introduces. Table 7 reports the decoding speed (Token per Second) with and without Spa on our machine. Using Spa will increase the inference time to 1.27 times that of the original model on average. In practice, we believe the additional overhead is negligible for developers. For example, Spa took 12.2 seconds to complete a HumanEval task, which is comparable to the original model’s 9.6 seconds.
There is negligible extra memory overhead because Spa only needs to buffer one additional logit tensor and a KV cache for a few masked tokens.
@@ Appendix G Hypothetical explanation for Attention dilution and Spa’s effectiveness
Spa is motivated by a recent study [23] and our empirical observations demonstrating the attention dilution issue. Our experiment results in Section 5 echo our observation and confirm the existence of attention dilution during code generation. Here we propose a detailed explanation for this phenomenon based on our knowledge and hypotheses. We believe it stems from two limitations in regular autoregressive decoding: (1) Distraction and (2) Error propagation.
@@
In contrast, the user prompt is persistently relevant throughout the generation since it represents the user’s intent. While self-generated tokens are also important context, they are less persistently related than task description in code generation. Amplifying the impact of the task description by Spa essentially enhances attention reliability, thereby mitigating distraction.
@@
Appendix H Setup of Large Language Models
Model Deployment. We downloaded and deployed LLMs from Huggingface. To expedite evaluations, we apply 8-bit quantization [13, 10] to all models. Prior studies [27, 20] have demonstrated that this approach has very little impact on LLM performance. All experiments were conducted on a 64-bit Ubuntu 22.04 LTS system, equipped with an AMD EPYC 7313 CPU, eight NVIDIA A5500 GPUs, and 512GB of memory. The experiments ran for approximately seven weeks.
Prompt Design. We use the original task descriptions from the datasets as prompts for the text-completion models, CodeLlama and CodeGen-Mono. For the three DeepSeek-Coder and StarCoder2 models, we format the prompts using the official chat template from HuggingFace.
@@ Appendix I Evaluating Spa on Other Generative Tasks
To evaluate the generalizability of Spa beyond code generation, we experiment Spa on other generative tasks, including TruthfulQA [28], GSM8K [8], MMLU [19], and BoolQ [7].
TruthfulQA [28] is a benchmark designed to measure models’ ability to avoid generating false or misleading information, requiring models to answer questions while remaining truthful. GSM8K [8] tests mathematical reasoning capabilities through grade school math word problems that require multi-step solutions. MMLU [19] evaluates models across 57 subjects, including elementary mathematics, US history, computer science, law, and more, comprehensively testing both breadth and depth of knowledge. It provides multiple-choice questions for LLMs to identify the correct answers. BoolQ [7] consists of naturally occurring yes/no questions from web queries, testing reading comprehension and binary classification abilities. We use Llama 3.1-Instruct-8B as our base model since the other models in our study are specifically fine-tuned for code tasks.
@@
In this section, we report a comprehensive evaluation of tuning Spa. To ensure fair comparisons, Spa is activated for all tasks in all experiments, unlike during inference time where Spa is only activated for failed tasks. Moreover, we tune Spa on all the tasks for better generalizability.
J.1 Analysis of Anchoring Strength
We study the relationship between code generation performance and the anchoring strength of Spa. Figure 6 illustrates the change in Pass@ for various values of across models and benchmarks ( represents the original model). We observe a roughly unimodal relationship between and performance: as increases, performance first improves, reaches an optimum, and then declines with further increases. We find the tuning is stable and transferrable. While the optimal varies slightly across different models and benchmarks, it tends to be model-dependent. Furthermore, we find serves as a universal value of anchoring strength that can improve performance across all cases.
J.2 Tuning Stability
@@
J.3 Cross-dataset & Cross-Model Tuning Evaluation
We investigate the transferability of this hyperparameter across different models and datasets. Firstly, we conduct a cross-dataset evaluation between HumanEval/HumanEval+ and MBPP/MBPP+, which have distinct prompt formats. We tune on HumanEval+ and evaluate Pass@ on MBPP and MBPP+, and vice versa888The “plus” versions of HumanEval and MBPP share identical prompts with their base counterparts, so we only tune once on the plus version. (denoted as Spacross-dataset). We calculate average Pass@ improvements on original and plus versions across all baseline models. Secondly, we perform a cross-model evaluation by tuning on one model and evaluating Pass@ on the remaining four. For each model, we compute the average Pass@ improvements across all the other models, for HumanEval/HumanEval+ and MBPP/MBPP+ respectively (denoted as Spacross-model). Similar to Section 5, Spa represents tuning within the split partial dataset, while Spa∗ represents tuning within the entire dataset.
@@
J.4 Comparision between Tuned Spa and Optimal Spa
@@
Table 13 reports optimal anchoring strength values . We observe the average value of can be used to effectively improve performance across all benchmarks for all LLMs.
@@
Unlike mainstream methods that optimize prompts, Spa enhances performance by optimizing model attention. We compare Spa with another attention steering method, PASTA [62]. PASTA’s implementation only supports LLAMA [48], LLAMA2 [49], and GPT-J [53]. We find that adapting PASTA to new models requires significant effort, so we only evaluate it on LLAMA-7B, LLAMA2-7B, and GPT-J-6B.
| HumanEval | HumanEval+ | Mbpp | Mbpp+ | BigCodeBench | |
| Pasta | +1.22 | +1.22 | +1.17 | +0.94 | +0.1 |
| Spa | +7.32 | +4.88 | +4.22 | +3.51 | +0.4 |
@@
Appendix M Performance Analysis by Different Generation Lengths
To evaluate how code length affects model performance, we analyze the length distribution of correctly and incorrectly generated code. We use the same set of experimental LLMs [36, 17, 41] as in Section 2 and experiment them on HumanEval [3]. We evaluate code length at three granularities: characters, words (tokens), and lines.
Figure 7 demonstrates that incorrectly generated code is significantly longer than correctly generated code across different granularity levels. This result echoes our findings about attention dilution—as the generated sequence grows longer, the model’s attention to the user prompt diminishes, leading to more errors.

### v5

@@ Appendix D Code Generation Examples
In the first example, CodeLlama (7B) overlooks the specified condition ”upper vowels.” In contrast, Spa enhances the model’s focus on the intended purpose. The code initializes all the upper vowels in the first line and correctly refers to it later.
@@ Appendix E Computational Cost of Spa
E.1 Latency
We analyzed the sources of inference overhead, which include running test cases, the two decoding processes, and the logit arithmetic operations. We observed that each attention augmentation in Spa results in an additional 0.6 times the base model’s latency, but this increase occurs only when test cases fail. Running test cases incurs only a very small cost (0.1s on average) compared to the total inference time (9.6s). Thus, only activating Spa when test failures are detected can reduce the overall overhead. The more accurate the base model, the less overhead Spa introduces. On average, Spa increases the inference time to 1.27 times that of the original model. We believe this overhead is acceptable in practical development.
@@ E.1 Latency
E.2 Memory
SPA requires the storage of the logit generated by the masked prompt embeddings. Theoretically, the additional memory requirement, denoted as , can be expressed as:
where represents the vocabulary size, is the token embedding dimension, and is the size of a single logit value.
Consider a vocabulary size , a token embedding dimension , and a logit size . Then the additional memory overhead is calculated as:
In practice, this memory overhead can be significantly reduced because many low-ranked tokens hardly contribute to the results. We can calculate the augmented logits by considering only the top-ranked tokens and ignoring the rest. For example, if we focus only on the top 100 logits, the overhead will dramatically decrease to .
E.3 Tradeoff Between Latency and Memory Usage via Parallelization
Currently, Spa is implemented by sequentially computing the logits. However, the latency can be further reduced by parallelizing the logits generation from the original and masked embeddings, as these computations are independent (Figure 1). This optimization, however, comes at the cost of increased memory usage (double VRAM or memory for forwarding). Given that LLM inference is memory-intensive, this introduces a tradeoff between reduced latency and higher memory consumption.
@@ Appendix G Hypothetical explanation for Attention dilution and Spa’s effectiveness
Spa is motivated by a recent study (Kou et al., 2024) and our empirical observations demonstrating the attention dilution issue. Our experiment results in Section 5 echo our observation and confirm the existence of attention dilution during code generation. Here we propose a detailed explanation for this phenomenon based on our knowledge and hypotheses. We believe it stems from two limitations in regular autoregressive decoding: (1) Distraction and (2) Error propagation.
@@
In contrast, the user prompt is persistently relevant throughout the generation since it represents the user’s intent. While self-generated tokens are also important context, they are less persistently related than task descriptions in code generation. Amplifying the impact of the task description by Spa essentially enhances attention reliability, thereby mitigating distraction.
@@
Appendix H Implementation and Deployment
H.1 Implementation of Spa
Spa is a model-agnostic algorithm and our implementation does not rely on specific models. All six models in our paper are built upon the Huggingface Transformer library, which offers APIs to directly access and edit token embeddings and logits. Particularly, the Spa generator inherits the Huggingface Transformers generation API.999https://huggingface.co/docs/transformers/en/main_classes/text_generation We leverage the hook to modify the logit calculation within the original generation pipeline. The API works for any LLM in the Huggingface model collections 101010https://huggingface.co/models with native hyperparameters such as temperature. We have released a PyPI library 111111https://pypi.org/project/anchoring/ for developers to quickly test Spa.
H.2 Model Deployment
We downloaded and deployed LLMs from Huggingface. To expedite evaluations, we apply 8-bit quantization (Frantar et al., 2023; Dettmers et al., 2022) to all models. Prior studies (Li et al., 2024; Huang et al., 2024) have demonstrated that this approach has very little impact on LLM performance. We set the Temperature to and the beam to for greedy decoding in all experiments, except for the one described in Appendix F. All experiments were conducted on a 64-bit Ubuntu 22.04 LTS system, equipped with an AMD EPYC 7313 CPU, eight NVIDIA A5500 GPUs, and 512GB of memory. The experiments ran for approximately seven weeks.
H.3 Prompt Design
We use the original task descriptions from the datasets as prompts for the text-completion models, CodeLlama and CodeGen-Mono. For the remaining models, we format the prompts using the official chat template from HuggingFace. All experiments are conducted in a zero-shot setting.
@@ Appendix I Evaluating Spa on Other Generative Tasks
To evaluate the generalizability of Spa beyond code generation, we experiment Spa on other generative tasks, including TruthfulQA (Lin et al., 2022), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021b), and BoolQ (Clark et al., 2019).
TruthfulQA (Lin et al., 2022) is a benchmark designed to measure models’ ability to avoid generating false or misleading information, requiring models to answer questions while remaining truthful. GSM8K (Cobbe et al., 2021) tests mathematical reasoning capabilities through grade school math word problems that require multi-step solutions. MMLU (Hendrycks et al., 2021b) evaluates models across 57 subjects, including elementary mathematics, US history, computer science, law, and more, comprehensively testing both breadth and depth of knowledge. It provides multiple-choice questions for LLMs to identify the correct answers. BoolQ (Clark et al., 2019) consists of naturally occurring yes/no questions from web queries, testing reading comprehension and binary classification abilities. We use Llama 3.1-Instruct-8B as our base model since the other models in our study are specifically fine-tuned for code tasks.
@@
As demonstrated in Section 5.6, different anchoring strengths lead to different performance and follow a simple unimodal pattern. In this section, we report more evaluation of tuning Spa. To ensure fair comparisons, Spa is activated for all tasks in all experiments, unlike during inference time where Spa is only activated for failed tasks. Moreover, we tune Spa on all the tasks for better generalizability.
J.1 Tuning Stability
@@
J.2 Cross-dataset & Cross-model Tuning Evaluation
We investigate the transferability of this hyperparameter across different models and datasets. Firstly, we conduct a cross-dataset evaluation between HumanEval/HumanEval+ and MBPP/MBPP+, which have distinct prompt formats. We tune on HumanEval+ and evaluate Pass@ on MBPP and MBPP+, and vice versa121212The “plus” versions of HumanEval and MBPP share identical prompts with their base counterparts, so we only tune once on the plus version. (denoted as Spacross-dataset). We calculate average Pass@ improvements on the original and plus versions across all baseline models. Secondly, we perform a cross-model evaluation by tuning on one model and evaluating Pass@ on the remaining four. For each model, we compute the average Pass@ improvements across all the other models, for HumanEval/HumanEval+ and MBPP/MBPP+ respectively (denoted as Spacross-model). Similar to Section 5, Spa represents tuning within the split partial dataset, while Spa∗ represents tuning within the entire dataset.
@@
We further investigate whether it is possible to find a universal anchoring strength that works for most scenarios. One potential value is the average of the tuned anchoring strengths across benchmarks for each model. We apply this fixed anchoring strength to all settings, denoted as Spa-preset. As shown in Table 13, although the generation accuracy decreases compared to using a tuned strength, Spa with the fixed strength can still outperform the baselines. It implies that Spa is effectively deployable in new scenarios once a reasonable value is set.
| Method | Pass@1 (%) | Time (Sec) |
| Base Model | 0 | 7.7 |
| Pasta | +1.2 | 48.8 |
| Self-Debugging | +4.2 | 27.3 |
| Self-Edit | +1.8 | 26.4 |
| Self-Planning | +3.6 | 21.6 |
| ReAct | +1.3 | 28.8 |
| Spa-fixed-strength | +5.0 | 9.8 |
| Spa | +7.7 | 9.8 |
J.3 Comparison between Tuned Spa and Optimal Spa
@@
Table 15 reports optimal anchoring strength values . We observe that the average value of can be used to effectively improve performance across all benchmarks for all LLMs.
@@
| HumanEval | HumanEval+ | Mbpp | Mbpp+ | BigCodeBench | |
| Pasta | +1.22 | +1.22 | +1.17 | +0.94 | +0.1 |
| Spa | +7.32 | +4.88 | +4.22 | +3.51 | +0.4 |
@@
Unlike mainstream methods that optimize prompts, Spa enhances performance by optimizing model attention. We compare Spa with another attention steering method, PASTA (Zhang et al., 2024). PASTA’s implementation only supports LLAMA (Touvron et al., 2023a), LLAMA2 (Touvron et al., 2023b), and GPT-J (Wang & Komatsuzaki, 2021). We find that adapting PASTA to new models requires significant effort, so we only evaluate it on LLAMA-7B, LLAMA2-7B, and GPT-J-6B.